If the ratio of two areas of circle A+B is 4, what is the ratio of teh radii of A+B
Answer by mananth(16946)

Area of circle A = pi*r1^2
Area of circle B = pi*r2^2

Area of circle A = pi*r1^2
---------------- -------
Area of circle B = pi*r2^2

4= r1^2/r2^2
r1/r2 = sqrt(4)
r1/r2= 2 (the positive value )